OpenStudy (anonymous):

SOLVE I5XI GREATER THEN = 0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\left| 5x \right|\ge0\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

\[-\infty<x<\infty\]